OpenStudy (anonymous):

A plane flew 1740 km in 6 hours with a tail wind. On the return trip, going into the wind, the plane flew 570 km in 3 hours. What is the plane's speed if there was no wind? What is the speed of the wind?

OpenStudy (anonymous):

Formula for Distance: R*T = D. Formula for Speed: D/T = R. Formula for Time: D/R = T. Hope this helps! :)